OpenStudy (anonymous):

Solve for the missing variable. Please show your steps to your solution. 6x+8=2x-28

OpenStudy (pinksapphire):

6x+8=2x-28 Combine like terms. -2x -2x 4x+8=-28 Combine like terms. -8 -8 4x/4=-36/4 Divide by 4 to get x by itself. x=9

OpenStudy (pinksapphire):

If you need any more help I can help you...

OpenStudy (anonymous):

Thank you